# Compose Your Policies! Improving Diffusion-based or Flow-based Robot Policies via Test-time Distribution-level Composition

### v1

@@ 4.1 Functional-level improvement
with strict inequality whenever the two models’ errors are not perfectly aligned.
@@ 4.2 System-level stability
where $F$ represents the underlying dynamics that map the score into the state update. Suppose $F$ is Lipschitz in $(x,s)$ with constants $L_{x}(t),L_{s}(t)$, and $\hat{s}$ is Lipschitz in $x$ with constant $\hat{\Lambda}(t)$. Assume the score error admits a uniform bound $\kappa(t)$. Define $\tilde{L}(t)=L_{x}(t)+L_{s}(t)\hat{\Lambda}(t)$. Then for all $T>0$,
@@ 4.3 Implications for policy composition
Corollary 4.1 (Convex score combination can reduce overall sampling error).
If a convex combination $s_{\mathrm{comp}}=ws_{1}+(1-w)s_{2}$ satisfies
@@
then
$\mathbb{E}\|x_{s_{\mathrm{comp}}}(T)-x^{*}(T)\|<\min_{i}\mathbb{E}\|x_{s_{i}}(T)-x^{*}(T)\|.$
@@ 5.1 Compositional Diffusion Models
1: Initialize noise trajectory $\tau_{N}\sim\mathcal{N}(0,I)$
2: for $w_{1}=0.0,0.1,\dots,0.9,1.0:$ // test-time searching
@@ 6.1 Experiment Settings
Environment Settings. We evaluate on Robomimic (Mandlekar et al., 2022), which includes three manipulation tasks (Can, Lift, Square), PushT (Florence et al., 2021), and RoboTwin (Mu et al., 2025), a suite of dual-arm collaborative tasks where we select representative ones from versions 1.0 and 2.0 (Chen et al., 2025). We further perform four real-world experiments: Place Bottles, Hang Mug, Close Table, and Punch Holes, with the setups in Fig. LABEL:fig:real_world_settings. More details are in App. H.
@@
Training and Testing Details. Since GPC is training-free, we directly use pre-trained policies trained based on their original settings (App. H). Each setting is evaluated over 200 rollouts (100 for RoboTwin), and we report the average success rate (SR). For composition, we employ our GCP and search over weighting coefficients from 0.0 to 1.0 in steps of 0.1.
@@ 7 Discussion
Limitations. Our GPC demonstrates clear effectiveness across a wide range of experiments. Despite this strength, certain limitations remain. First, test-time weight search is restricted by a fixed discretization, which may overlook optimal values; future work could explore adaptive or automatic search strategies. Second, we mainly study dual-policy composition, while scaling to more policies increases computation. Addressing this may require feature sharing or compact representations to enable efficient multi-policy integration.
Conclusion. We introduced General Policy Composition, a training-free framework that improves robotic control by combining the distributional scores of pre-trained policies. Our theoretical analysis establishes that convex score composition leads to step-wise and trajectory-level improvements, while our experiments on diverse benchmarks and real-world setups confirm consistent performance gains. GPC is simple, versatile, and widely applicable, providing a foundation for future research in policy composition as a means to enhance control performance without additional training resources.
@@ Appendix A Assumptions, Notation, and Preliminary Facts
$\mathbb{E}\|x_{\hat{s}}(T)-x^{*}(T)\|\;\leq\;\Bigg(\int_{0}^{T}e^{\,2\int_{t}^{T}\tilde{L}(\tau)\,d\tau}\,L_{s}(t)^{2}\,dt\Bigg)^{\!1/2}\Bigg(\int_{0}^{T}\kappa(t)^{2}\,dt\Bigg)^{\!1/2}.$ (13)
Please see the detailed proof as follows.
@@ D. Expectation and a Readable Upper Bound
$\mathbb{E}\|e(T)\|\;\leq\;\Bigg(\int_{0}^{T}e^{\,2\int_{t}^{T}\tilde{L}(\tau)\,d\tau}\,L_{s}(t)^{2}\,dt\Bigg)^{\!1/2}\Bigg(\int_{0}^{T}\mathbb{E}\|\Delta_{s}(t,x^{*}(t))\|^{2}\,dt\Bigg)^{\!1/2}$
@@
Let $s_{\mathrm{comp}}=ws_{1}+(1-w)s_{2}$ with $w\in(0,1)$. If
$\int_{0}^{T}\mathbb{E}\|s_{\mathrm{comp}}-s^{*}\|^{2}\,dt\;<\;\min_{i\in\{1,2\}}\int_{0}^{T}\mathbb{E}\|s_{i}-s^{*}\|^{2}\,dt,$
then
$\mathbb{E}\|x_{s_{\mathrm{comp}}}(T)-x^{*}(T)\|\;<\;\min_{i\in\{1,2\}}\mathbb{E}\|x_{s_{i}}(T)-x^{*}(T)\|.$
@@
By Proposition 4.2, for any score estimator $\hat{s}$ we have
$\mathbb{E}\|x_{\hat{s}}(T)-x^{*}(T)\|\;\leq\;\exp\!\Big(\!\int_{0}^{T}\tilde{L}\Big)\,\Big(\int_{0}^{T}L_{s}(t)^{2}\,dt\Big)^{\!1/2}\,\Big(\int_{0}^{T}\mathbb{E}\|\hat{s}-s^{*}\|^{2}dt\Big)^{\!1/2},$
for all choices of $\hat{s}$. Applying this inequality to $\hat{s}=s_{\mathrm{comp}}$ and to $\hat{s}=s_{i}$ $(i=1,2)$, and and noting that Proposition 4.1 guarantees the existence of a convex weight $w^{*}\in(0,1)$ for which the integrated MSE of $s_{\mathrm{comp}}$ is strictly smaller than that of each $s_{i}$,
$\int_{0}^{T}\mathbb{E}\|s_{\mathrm{comp}}-s^{*}\|^{2}\,dt\;<\;\min_{i\in\{1,2\}}\int_{0}^{T}\mathbb{E}\|s_{i}-s^{*}\|^{2}\,dt,$
we conclude that the corresponding terminal trajectory error of $s_{\mathrm{comp}}$ is also strictly smaller than that of either base estimator. ∎
@@ Appendix G The Flexibility of GPC with Any Prediction Types
Appendix H Experiment Details
H.1 Robomimic
@@ H.3 RoboTwin
H.4 Real-world Experiments
[…]
@@ H.4 Real-world Experiments
H.5 Notation for GPC Flexibility
@@
Appendix I Additional Experimental Results
@@
Appendix J Visualization on Robot Tasks
@@
Appendix K Detailed Literature Review
@@
K.1 Compositional Generative Modeling
@@
K.2 Diffusion Models in Robot Learning
@@
K.3 Non-Diffusion-based Models in Robot Learning
@@
Appendix L Future Work

### v2

@@ 4.1 Functional-level improvement
with strict inequality in most non-trivial cases.
@@ 4.2 System-level stability
where $F$ represents the underlying dynamics that map the score into the state update. Suppose $F$ is Lipschitz in $(x,s)$ with constants $L_{x}(t),L_{s}(t)$, and $\hat{s}$ is Lipschitz in $x$ with constant $\hat{\Lambda}(t)$. Assume the score error admits a uniform bound $\kappa(t)$ (i.e., $\|\hat{s}-s^{*}\|\leq\kappa(t)$). Define $\tilde{L}(t)=L_{x}(t)+L_{s}(t)\hat{\Lambda}(t)$. Then for all $T>0$,
@@ 4.3 Implications for policy composition
Corollary 4.1 (Convex score combination tightens the sampling error bound).
[…]
@@
then the corresponding theoretical error bound is strictly reduced:
$\mathcal{B}(s_{\mathrm{comp}})<\min_{i}\mathcal{B}(s_{i}).$
@@ 5.1 Compositional Diffusion Models
1: for $w_{1}=0.0,0.1,\dots,0.9,1.0:$ // test-time searching
2: Initialize noise trajectory $\tau_{N}\sim\mathcal{N}(0,I)$ per action
@@ 6.1 Experiment Settings
Environment Settings. We evaluate on Robomimic (Mandlekar et al., 2022), which includes three manipulation tasks (Can, Lift, Square), PushT (Florence et al., 2021), and RoboTwin (Mu et al., 2025), a suite of dual-arm collaborative tasks where we select representative ones from versions 1.0 and 2.0 (Chen et al., 2025). We further perform four real-world experiments: Place Bottles, Hang Mug, Close Table, and Punch Holes, with the setups in Fig. LABEL:fig:real_world_settings. More details are in App. I.
@@
Training and Testing Details. Since GPC is training-free, we directly use pre-trained policies trained based on their original settings (more details are in App. I). Each setting is evaluated over 200 rollouts (100 for RoboTwin), and we report the average success rate (SR). For composition, we employ our GCP and search over weighting coefficients from 0.0 to 1.0 in steps of 0.1.
Optimized weight searching (Optional). Although GPC in principle allows arbitrary convex weights, our framework naturally suggests that the policy with the stronger score should receive a larger weight. This intuition is empirically validated by our findings in Sec. 6.3. Therefore, in practical use, once the relative strengths of the base policies are known, one can directly bias the stronger policy to have weight >>$0.5$, instead of exhaustively searching over the entire interval $[0.1,0.9]$. This simple heuristic both respects the underlying score-composition principle and significantly reduces the searching time.
@@ 6.3 Influence of Weight Configurations on GPC Performance
Table 6: Comparison of training/finetuning time vs. GPC weight search. $T_{\text{eval}}=N_{\text{rollout}}\times T_{\text{per\_rollout}}$. For RoboMimic, $T_{\text{eval}}^{\text{RoboMimic}}$$\approx 200*5\text{s}=0.27\text{ hr}$. For real-world, $T_{\text{eval}}^{\text{Real}}$$\approx 20*30\text{s}=0.17\text{ hr}$.
Method Setting Time cost
Training from scratch 1M+ demos, N GPUs 14d (OpenVLA 7B), 30d (RDT 1B)
Finetuning 100 demos, 1 GPU >>5h (DP 200M+), >>8h (RDT 1B)
GPC (full search) 9 weights (w=0.1:0.9w=0.1{:}0.9) $9\,T_{\text{eval}}$ ($\sim$ 2.5hrs)
GPC (optimized) 4 weights ($w_{\text{strong}}\in[0.6,0.9]$) $4\,T_{\text{eval}}$ ($\sim$ 1hr)
Table 7: Per–action-chunk inference latency in RoboMimic. The overhead of GPC is modest and purely computational.
Method Time per chunk (s)
DP 0.09
Florence-Policy-D 0.06
GPC (DP + FP-D) 0.13
6.4 Time and computational efficiency of GPC
GPC introduces two main sources of extra cost compared to using a single base policy: (i) additional weight search evaluations (Snell et al., 2024), and (ii) increased inference cost per action chunk.
Weight search as repeated evaluations. During weight search, we fix a candidate weight configuration, run $N$ rollouts to estimate the SR, and then move to the next configuration. Thus, the total search cost is $T_{\text{search}}=N_{\text{search}}*T_{\text{eval}}$, $\quad T_{\text{eval}}=N_{\text{rollout}}*T_{\text{per\_rollout}}.$ As shown in Tab. 7, sweeping $w$ from $0.1$ to $0.9$ leads to $N_{\text{search}}=9$ and about $2.5$ hours. With our optimized searching method, we can restrict the search to $w_{\text{strong}}\in[0.6,0.9]$, reducing to $N_{\text{cfg}}=4$ ($\sim 1$ hour). Compared to days of training from scratch or hours of finetuning, this makes GPC a highly efficient alternative.
Inference-time overhead. During rollout, GPC incurs extra compute from querying multiple base policies per action chunk. Tab. 7 reports the measured inference latency in RoboMimic: the GPC composition increases latency from $0.09$ s to $0.13$ s per chunk, which is modest in practice. This overhead is purely computational and can be further reduced with stronger hardware, optimized inference runtimes, or future engineering improvements (e.g., model compression (Høeg et al., 2024) or distillation (Liu et al., 2026)).
@@ 6.5 Comprehensive Analysis of GPC Effectiveness
GPC with heterogeneous inference steps and chunk sizes. GPC is also flexible with different inference steps and chunk sizes. For diffusion steps, we simply pick a unified sampler and number of inference steps at test time and run all policies under this common configuration, e.g., the results in RoboTwin with DP+RDT. For action-chunk mismatch (assume $H_{A}\geq H_{B}$), we sample a shared noise trajectory of length $H_{A}$, apply policy B only on the first $H_{B}$ steps, and take a convex combination of scores on the overlap while keeping the tail from policy A. Results in Tab. 9 confirm that this heterogeneous-chunk composition also yields clear gains.
GPC with multiple base policies. The score-composition rule of GPC directly extends to three or more base policies via a convex combination of their scores. Tab. 9 reports results with three policies on RoboMimic. Three-policy GPC either matches or improves upon the best two-policy configuration, and consistently outperforms the individual base policies. This shows that GPC scales beyond pairwise composition and can effectively leverage diversity across multiple pretrained policies.
Table 8: GPC can be applied with different action-chunk lengths and infer time steps. Results are in Robomimic.
Method Setting Success Rate
DP DDPM, chunk 8, 5 steps 0.50
Florence-Policy-D DDIM, chunk 16, 10 steps 0.53
GPC (DP+FP-D) DDIM, chunk 16, 10 steps 0.66
Table 9: GPC with three base policies on RoboMimic.
Method Can Lift Square
Flow Policy 0.95 0.13 0.77
Florence-Policy-F 0.89 0.98 0.88
$\pi_{0}$ 0.61 0.96 0.92
GPC (best 2-policy) 0.99 1.00 0.94
[…]
@@ 7 Discussion
Limitations. Our GPC demonstrates clear effectiveness across a wide range of experiments. Despite this strength, certain limitations remain. First, test-time weight search is restricted by a fixed discretization, which may overlook optimal values; future work could explore adaptive or automatic search strategies. Second, we mainly study dual/triple-policy composition, while scaling to more policies increases computation. Addressing this may require feature sharing or compact representations to enable efficient multi-policy integration.
Conclusion. We introduced General Policy Composition, a training-free framework that improves robotic control by combining the distributional scores of pre-trained policies. Our theoretical analysis establishes that convex score composition leads to step-wise and trajectory-level improvements, while our experiments on diverse benchmarks and real-world setups confirm consistent performance gains. GPC is simple, versatile, and widely applicable, providing a foundation for future research in policy composition as a means to enhance performance without additional training resources.
@@ Appendix A Assumptions, Notation, and Preliminary Facts
Remark on strict improvement.
Strict improvement requires the unconstrained optimum to lie within $(0,1)$. If one estimator significantly outperforms the other, the optimal weight lies on the boundary, recovering the best single-estimator performance.
@@
$\mathbb{E}\|x_{\hat{s}}(T)-x^{*}(T)\|\;\leq\;\Bigg(\int_{0}^{T}e^{\,2\int_{t}^{T}\tilde{L}(\tau)\,d\tau}\,L_{s}(t)^{2}\,dt\Bigg)^{\!1/2}\Bigg(\int_{0}^{T}\kappa(t)^{2}\,dt\Bigg)^{\!1/2},$ (13)
where $\kappa(t)$ is a nonnegative function $\kappa(t)$ such that for all $x$,
$\|\Delta_{s}(t,x)\|\leq\kappa(t).$
(i.e., Score error bound assumption). Please see the detailed proof as follows.
@@ D. Expectation and a Readable Upper Bound
$\mathbb{E}\|e(T)\|\;\leq\;\Bigg(\int_{0}^{T}e^{\,2\int_{t}^{T}\tilde{L}(\tau)\,d\tau}\,L_{s}(t)^{2}\,dt\Bigg)^{\!1/2}\Bigg(\int_{0}^{T}\mathbb{E}\|\Delta_{s}(t,x^{*}(t))\|^{2}\,dt\Bigg)^{\!1/2}$ (21)
@@
[…]
then the corresponding theoretical error bound is strictly reduced:
$\mathcal{B}(s_{\mathrm{comp}})<\min_{i}\mathcal{B}(s_{i}).$
@@
From Proposition 4.2, the expected trajectory error for any estimator $\hat{s}$ is bounded by:
$\mathbb{E}\|x_{\hat{s}}(T)-x^{*}(T)\|\;\leq\;\mathcal{B}(\hat{s}):=\Bigg(\int_{0}^{T}e^{\,2\int_{t}^{T}\tilde{L}(\tau)\,d\tau}\,L_{s}(t)^{2}\,dt\Bigg)^{\!1/2}\cdot\left(\int_{0}^{T}\mathbb{E}\|\hat{s}-s^{*}\|^{2}dt\right)^{1/2},$
where $\Bigg(\int_{0}^{T}e^{\,2\int_{t}^{T}\tilde{L}(\tau)\,d\tau}\,L_{s}(t)^{2}\,dt\Bigg)^{\!1/2}>0$ is a system-dependent constant independent of the score estimator. The bound function $\mathcal{B}(\cdot)$ is strictly increasing with respect to the integrated mean-squared score error (MSE). Since the premise states that the integrated MSE of the composite score $s_{\mathrm{comp}}$ is strictly smaller than that of the individual estimators $s_{i}$, it follows immediately that:
$\mathcal{B}(s_{\mathrm{comp}})<\min_{i}\mathcal{B}(s_{i}).$
Thus, convex score composition strictly tightens the theoretical guarantee on the trajectory simulation error. ∎
@@ Appendix G The Flexibility of GPC with Any Prediction Types
Appendix H In-depth analysis on what leads to the success of GPC
Intuitively, GPC can be viewed as forming a product-of-experts distribution: the composed score is a convex combination of individual scores, corresponding to a (re-weighted) product of their probability densities. A higher weight on one policy simply means its density contributes more strongly to the final target distribution, concentrating probability mass on trajectories that both experts consider likely.
H.1 Theoretical perspective: why convex score composition can be better than individual policies
Based on Prop. 4.1, we show that at each diffusion timestep there exists a convex weight $w^{*}$ such that the composed score has a smaller error with respect to the ideal score $s^{*}$ than any individual policy. Prop. 4.2 extends this from a single step to the entire denoising trajectory: if at each step the composed score is closer to $s^{*}$, then the error along the whole trajectory accumulates more favorably, leading to a trajectory distribution closer to the ideal one. In other words, once per-step improvement in score quality is established, the stability of the generative process allows this advantage to propagate along the full trajectory, making convex score composition provably superior to relying on a single policy.
H.2 Empirical perspective: what drives larger gains for some combinations
At a high level, we view GPC as a way to aggregate the “good knowledge” learned by different base policies into a single, higher-likelihood target distribution, so that sampling from the composed score produces higher-quality trajectories than sampling from any individual policy alone. In practice, what counts as “good knowledge” is shaped by several factors:
Complementary modalities provide richer information. When GPC combines policies trained on different modalities (e.g., RGB vs. point cloud), the composed score effectively leverages complementary views of the scene: RGB captures appearance, texture, and color cues, while point clouds provide precise 3D geometry and depth structure. This multimodal fusion reduces perceptual ambiguity compared to using either modality alone, and can be viewed as increasing the “informational richness” available to the composed policy. The DP+DP3 combination is a good example of this, and the visualizations of their sample distributions in Fig. 5(a) support this interpretation.
Diverse architectures capture different inductive biases. Even when policies share the same input modality, different architectures can encode different inductive biases and error patterns. For instance, $\pi 0$ is DiT-based, whereas Flow Policy uses a U-Net-style backbone. They may model similar underlying action distributions but emphasize different structures in the data. GPC can exploit this diversity by aggregating their strengths while averaging out idiosyncratic weaknesses, which we see reflected in improved performance and the qualitative analysis in Fig. 5(b).
Task-specific strengths shape composition benefits. Task choice also matters: in some tasks, one base policy may be significantly weaker. In such cases, naive averaging (or an unbalanced weight) can let the weaker policy drag down the overall performance. Our analysis in Sec. 6.3 (Finding 3) shows that GPC performs best when the *better-performing base policy receives a larger weight*, which is especially important when task difficulty or mismatch affects one policy more than the other. Thus, task-specific policy strengths and the ability to adjust weights are key factors behind the observed gains.
Weight selection matters in the composed distribution. Our theory guarantees the existence of an optimal weight $w^{*}$, but in practice GPC still requires choosing concrete weights. These weights control how much each base distribution contributes to the final composed distribution and therefore directly influence how much probability mass is placed on high-quality trajectories. This is why we performed extensive weight-sweep studies: to show how different weights affect performance, and to provide empirical guidance on weight selection for real deployments. We view designing better, possibly adaptive, strategies to approximate $w^{*}$ as an important direction for future work.
Appendix I Experiment Details
I.1 Robomimic
@@ I.3 RoboTwin
Task Prompt in RoboTwin. Here we present the detailed text description and its corresponding text prompt for VLAs. For example, for the Place Burger task, the task description and schema are:
  • Full description: “Use dual arm to pick the hamburg and frenchfries and put them onto the tray.”
  • Schema: “A denotes the hamburg, B denotes the tray, C denotes the frenchfries”
During training, the VLA receives diverse paraphrased prompts for this task, such as:
  • “Use both arms to move A and C to B.”
  • “Lift A and C, placing them neatly on B”
At test time, we use unseen paraphrases with the same schema, e.g..:
  • “Pick A and C, then place them on B.”
  • “Grab A and C together, setting them on B.”
@@
I.4 Real-world Experiments
[…]
I.5 Notation for GPC Flexibility
@@
Appendix J Additional Experimental Results
@@
Appendix K Visualization on Robot Tasks
@@
Appendix L Detailed Literature Review
@@
L.1 Compositional Generative Modeling
@@
L.2 Diffusion Models in Robot Learning
@@
L.3 Non-Diffusion-based Models in Robot Learning
@@
Appendix M Future Work